Find the constants C0, C1, and x1 so that the quadrature formulae z0^1 f(x) dx = C0 f(0) + C1 f(x1),





Has the highest possible degree of precision

1
Expert's answer
2022-03-10T15:16:29-0500

Due to the fact that is the boundary point of the segment, the highest possible degree of precision is "2" (number of nodes minus "2"). Solve the equation for "f(x)=ax^2+bx+c" :

"\int_0^1ax^2+bx+c=c_0f(0)+c_1f(x_1)\\
\frac{a}{3}+\frac{b}{2}+c=c_0c+c_1ax_1^2+c_1bx_1+c_1x"

Due to the fact that "a, b, c" are arbitrary numbers, the following system of equations is obtained (grouping the terms by "a, b, c") :

"\begin{cases}
\frac{1}{3}=c_1x_1^2\\
\frac{1}{2}=c_1x_1\\
1=c_0+c_1
\end{cases}"

Solving the system we get the answer: "x_1=\frac{2}{3},\ c_1=\frac{3}{4},\ c_0=\frac{1}{4}"